# Redefining Developer Assistance: Through Large Language Models in Software Ecosystem

### 09 Dec 2023 v1

@@ Abstract.
copyright: none
1. Introduction
The advent of Large Language Models (LLMs) has brought a revolutionary shift in various industries, with a significant impact in the field of software development. These models, ranging from General-Purpose Transformers like GPT by OpenAI (Brown et al., 2020) to specialized ones such as LLaMA (Introducing LLaMA: A foundational, [n. d.]) and Alpaca (Taori et al., 2023), have redefined human-computer interaction and automation in software engineering tasks. In the software industry, LLMs have been instrumental in automating tasks such as code generation (Ross et al., 2023; Zhang et al., 2023), test case creation (Schäfer et al., 2023), and program debugging (Chen et al., 2023). Tools like OpenAI’s Codex (che, 2021) and GitHub Copilot (Dakhel et al., 2023) are exemplary in facilitating code-related tasks and enhancing developer productivity.
However, despite these advancements in the field, the software industry  [^1] https://insights.sei.cmu.edu/blog/application-of-large-language-models-llms-in-software-engineering-overblown-hype-or-disruptive-change/ confronts substantial challenges beyond the realm of coding. While coding is an integral part of a developer’s role, it occupies about 59.5% of their work time  [^2] https://www.software.com/reports/future-of-work  [^3] https://thenewstack.io/how-much-time-do-developers-spend-actually-writing-code/, reflecting a notable focus on this aspect of software development. Surprisingly, a significant portion of a developer’s time, nearly 40-45%, involves tasks beyond coding. These include code maintenance  [^4] https://thenewstack.io/how-much-time-do-developers-spend-actually-writing-code/, testing, security issues  [^5] https://www.techtarget.com/searchitoperations/feature/The-promises-and-risks-of-AI-in-software-development, and other non-coding responsibilities that are essential for the holistic development of software. This division of time highlights the necessity for tools that cater not only to the coding aspect but also to the broader scope of software development. Tasks involving key concepts from Named Entity Recognition (NER), Relation Extraction (RE), Link Prediction (LP), Question Answering (QA), and Forum Answer Ranking (FAR) are pivotal in this broader scope. They play a critical role in interpreting complex technical documentation  [^6] https://pub.aimind.so/interrogate-your-technical-documentation-using-free-and-paid-llms-f2a7664ff2bd, managing software project requirements  [^7] https://www.linkedin.com/pulse/requirement-engineering-using-llm-madhavan-vivekanandan/, and facilitating effective communication within developer communities  [^8] https://www.smehorizon.com/large-language-models-5-takeaways-for-smes/.
@@ 1. Introduction
Figure 1. Distribution of DevAssistLlama dataset
@@
Table 1. Quantitative Analysis of NLP Models on Named Entity Recognition, Relation Extraction, and Link Prediction.
@@
Table 2. FAR Comparison of Different Models
@@
Table 3. QA Performance of Different Models
2. Dataset
@@
Table 4. Dataset Statistics
@@
Named entity recognition (NER) (Anonymous, 2023; Tabassum et al., 2020; Nguyen et al., 2023): NER is an elementary method for automated requirement extraction (Haris and Kurniawan, 2021), code document analysis (Chew et al., 2023) and bug tracking and resolution. The task is to extract the entities of specific entity types given a software related textual information. We consider four software system related NER datasets. The  DistalaNER (Anonymous, 2023) data consists of nine different software entity types. The textual information is the bug description. The entities are packages, operating system, organization, commands, error, file extension, peripheral, software component and computer architecture. Total number of instances present in the dataset is 170K. We used the  Stackoverflow and  Github data proposed in paper (Tabassum et al., 2020). In these datasets, the texts are the sentences of the questions. Both of them has around  20 entity types where 8 code related entities and 12 natural language entities. The stackoverflow dataset consists of 15,372 sentences with 19,122 entities. The github data consists of 6,510 sentences with 10,963 entities. The Wikiser (Nguyen et al., 2023) dataset consists of 12 entity types. It contains the wikipedia textual information. This dataset consists of 1.7M sentences with 79K unique software entities.
Relation extraction (RE) (Anonymous, 2023): In this dataset, textual information, along with two entities, serves as input, and the relation type acts as the output. It encompasses five relation types: dependency, conflict, affected version, cause and effect, and interaction/control. The textual information comprises bugs from the Ubuntu ecosystem. This dataset contains 500 instances.
Forum answer ranking (FAR) (Abdelaziz et al., 2022): We use the forum answer ranking dataset from the Blanca benchmark proposed in paper (Abdelaziz et al., 2022). The task is to predict the best answers relative to other answers provided for a given question in a forum. The dataset consists of 500K questions, each having at least 3 answers, with an average of  4.9 answers per question.
Forum link prediction (LP) (Abdelaziz et al., 2022): The task involves predicting whether two questions are related or not. The dataset consists of 23,516 pairs of posts, with 11,758 related pairs and 11,758 unrelated pairs.
@@
3. Developer assistance through instruction tuning
In this work, our main objective is to build a model which would assist developers to address natural language software queries in various stages. We utilize the instruction tuning method on pretrained LLMs to achieve this objective. Instruction tuning is one of the fine tuning methods for aligning the llm to specific tasks or specific domain. For each pair of instruction and desired response $(x,y)$ in the instruction dataset, we update the model parameters $\theta$ to reduce the difference between the model’s output $y^{{}^{\prime}}$ and desired output $y$. Further, this process iteratively update $\theta$ to improve the model’s performance in following instructions and producing the desired outputs. In our work, this methodology involves the refinement of pre-existing llama2-13b. There are various existing works on instruction tuning where the main focus is to solve specific tasks (Zhou et al., 2023; Lu et al., 2023). Whereas we investigate how these pretrained LLMs as well as instruction tuned LLM can be employed as a component in the pipeline of various software system tasks. In the subsequent sections, we describe the construction of instruction dataset, instruction tuning process and evaluation tasks.
@@ 3. Developer assistance through instruction tuning
Assume that you are an answer ranker in a community question answering portal. Given the question, rank the following answers based on their relevance from most relevant (1) to least relevant (0). If an answer is completely irrelevant, it should be ranked last. Strictly follow the Output Format: ”OPTION [answer number]: Rank [ranking number]”. DO NOT include any other additional text.
Table 5. System Tasks and Instruction Prompts
3.1. Instruction dataset construction
@@ 3.3. Negative sampling
4. Experimental setup
4.1. Model setup
In our study, we employ a model setup inspired by (Chiang et al., 2023). The cornerstone of our experimentation is Llama 2, fine-tuned to enhance its efficacy in our designated tasks. This approach is in line with cutting-edge methodologies in natural language processing. For evaluation, we select up to 1,000 instances from three distinct random splits, ensuring a diverse and representative test dataset. This strategy aims to emulate real-world scenarios and offers a comprehensive evaluation of the model’s capabilities.
4.2. Compared models
We evaluate the performance of several prominent Large language models. ChatGPT (gpt-3.5-turbo-0301) is renowned for its conversational abilities and versatility in various language tasks, initially based on the GPT-2 model from (Radford et al., 2019). Vicuna, credited to (Chiang et al., 2023), stands out for its efficiency, being fine-tuned on a LLaMA base model using conversational data. Similarly, Alpaca, developed by (Taori et al., 2023), uses the LLaMA architecture but is specifically tuned for domain-specific dialogues. Our focus, Llama 2, is fine-tuned following (tou, 2023) guidelines, showcasing its capabilities in complex language processing scenarios and offering insights into the evolving landscape of vaious NLP tasks.
4.3. LoRA approach for fine-tuning
Low-Rank Adapters (LoRA) represent a pivotal advancement in the field of natural language processing, particularly in the fine-tuning of large language models. First introduced by (Hu et al., 2021), LoRA allows for the efficient and effective modification of pre-trained models, significantly reducing the computational overhead typically associated with training large-scale neural networks. In our implementation, we integrate LoRA adapters directly into the attention mechanisms and train both the language model (LM) heads and the adapters concurrently, diverging from traditional methods that prioritize the initial training of embeddings alone, as seen in models like the Chinese LLaMA (Cui et al., 2023). Our infrastructure utilizes the robust capabilities of six NVIDIA A100 GPUs, each with 46GB of memory, enabling us to train our model, DevAssistLlama, over three epochs spanning a total of 63 hours. We showcased important hyperparameters in Table 6.
5. Results
@@ 5.1. Evaluation metrics
Learning Rate: 2e-4— Max sequence length: 4096—
LoRA Alpha: 16— Training Precision: FP16—
Epochs: 3— Optim: paged_adamw_32bit—
fp16=True— bf16=True— Warmup ratio=0.03—
LR scheduler=Cosine— Trainer= SFTT—
@@
Table 6. Summary of important hyperparameters
6. Limitation and future work

### 09 Mar 2024 v2

@@ Abstract
1 Introduction
The advent of Large Language Models (LLMs) has brought a revolutionary shift in various industries, with a significant impact in the field of software development. These models, ranging from General-Purpose Transformers like GPT by OpenAI Brown et al. (2020) to specialized ones such as LLaMA Introducing LLaMA: A foundational and Alpaca Taori et al. (2023), have redefined human-computer interaction and automation in software engineering tasks. In the software industry, LLMs have been instrumental in automating tasks such as code generation Ross et al. (2023); Zhang et al. (2023), test case creation Schäfer et al. (2023), and program debugging Chen et al. (2023). Tools like OpenAI’s Codex che (2021) and GitHub Copilot Dakhel et al. (2023) are exemplary in facilitating code-related tasks and enhancing developer productivity.
However, despite these advancements in the field, the software industry  [^1] https://insights.sei.cmu.edu/blog/application-of-large-language-models-llms-in-software-engineering-overblown-hype-or-disruptive-change/ confronts substantial challenges beyond the realm of coding. While coding is an integral part of a developer’s role, it occupies about 59.5% of their work time  [^2] https://www.software.com/reports/future-of-work  [^3] https://thenewstack.io/how-much-time-do-developers-spend-actually-writing-code/, reflecting a notable focus on this aspect of software development. Surprisingly, a significant portion of a developer’s time, nearly 40-45%, involves tasks beyond coding. These include code maintenance  [^4] https://thenewstack.io/how-much-time-do-developers-spend-actually-writing-code/, testing, security issues  [^5] https://www.techtarget.com/searchitoperations/feature/The-promises-and-risks-of-AI-in-software-development, and other non-coding responsibilities that are essential for the holistic development of software. This division of time highlights the necessity for tools that cater not only to the coding aspect but also to the broader scope of software development. Tasks involving key concepts from Named Entity Recognition (NER), Relation Extraction (RE), Link Prediction (LP), Question Answering (QA), and Forum Answer Ranking (FAR) are pivotal in this broader scope Banerjee et al. (2024b, a); Hazra et al. (2023, 2024). They play a critical role in interpreting complex technical documentation  [^6] https://pub.aimind.so/interrogate-your-technical-documentation-using-free-and-paid-llms-f2a7664ff2bd, managing software project requirements  [^7] https://www.linkedin.com/pulse/requirement-engineering-using-llm-madhavan-vivekanandan/, and facilitating effective communication within developer communities  [^8] https://www.smehorizon.com/large-language-models-5-takeaways-for-smes/.
@@ 1 Introduction
Figure 1: Distribution of DevAssistLlama dataset
@@
Table 1: Quantitative Analysis of NLP Models on Named Entity Recognition, Relation Extraction, and Link Prediction.
@@
Table 2: FAR Comparison of Different Models
@@
Table 3: QA Performance of Different Models
2 Dataset
@@
Table 4: Dataset Statistics
@@
Named entity recognition (NER) Anonymous (2023); Tabassum et al. (2020); Nguyen et al. (2023): NER is an elementary method for automated requirement extraction Haris and Kurniawan (2021), code document analysis Chew et al. (2023) and bug tracking and resolution. The task is to extract the entities of specific entity types given a software related textual information. We consider four software system related NER datasets. The  DistalaNER Anonymous (2023) data consists of nine different software entity types. The textual information is the bug description. The entities are packages, operating system, organization, commands, error, file extension, peripheral, software component and computer architecture. Total number of instances present in the dataset is 170K. We used the  Stackoverflow and  Github data proposed in paper Tabassum et al. (2020). In these datasets, the texts are the sentences of the questions. Both of them has around  20 entity types where 8 code related entities and 12 natural language entities. The stackoverflow dataset consists of 15,372 sentences with 19,122 entities. The github data consists of 6,510 sentences with 10,963 entities. The Wikiser Nguyen et al. (2023) dataset consists of 12 entity types. It contains the wikipedia textual information. This dataset consists of 1.7M sentences with 79K unique software entities.
Relation extraction (RE) Anonymous (2023): In this dataset, textual information, along with two entities, serves as input, and the relation type acts as the output. It encompasses five relation types: dependency, conflict, affected version, cause and effect, and interaction/control. The textual information comprises bugs from the Ubuntu ecosystem. This dataset contains 500 instances.
Forum answer ranking (FAR) Abdelaziz et al. (2022): We use the forum answer ranking dataset from the Blanca benchmark proposed in paper Abdelaziz et al. (2022). The task is to predict the best answers relative to other answers provided for a given question in a forum. The dataset consists of 500K questions, each having at least 3 answers, with an average of  4.9 answers per question.
Forum link prediction (LP) Abdelaziz et al. (2022): The task involves predicting whether two questions are related or not. The dataset consists of 23,516 pairs of posts, with 11,758 related pairs and 11,758 unrelated pairs.
@@
3 Developer assistance through instruction tuning
In this work, our main objective is to build a model which would assist developers to address natural language software queries in various stages. We utilize the instruction tuning method on pretrained LLMs to achieve this objective. Instruction tuning is one of the fine tuning methods for aligning the llm to specific tasks or specific domain. For each pair of instruction and desired response $(x,y)$ in the instruction dataset, we update the model parameters $\theta$ to reduce the difference between the model’s output $y^{{}^{\prime}}$ and desired output $y$. Further, this process iteratively update $\theta$ to improve the model’s performance in following instructions and producing the desired outputs. In our work, this methodology involves the refinement of pre-existing llama2-13b. There are various existing works on instruction tuning where the main focus is to solve specific tasks Zhou et al. (2023); Lu et al. (2023). Whereas we investigate how these pretrained LLMs as well as instruction tuned LLM can be employed as a component in the pipeline of various software system tasks. In the subsequent sections, we describe the construction of instruction dataset, instruction tuning process and evaluation tasks.
@@ 3 Developer assistance through instruction tuning
Assume that you are an answer ranker in a community question answering portal. Given the question, rank the following answers based on their relevance from most relevant (1) to least relevant (0). If an answer is completely irrelevant, it should be ranked last. Strictly follow the Output Format: "OPTION [answer number]: Rank [ranking number]". DO NOT include any other additional text.
Table 5: System Tasks and Instruction Prompts
3.1 Instruction dataset construction
@@ 3.3 Negative sampling
4 Experimental setup
4.1 Model setup
In our study, we employ a model setup inspired by Chiang et al. (2023). The cornerstone of our experimentation is Llama 2, fine-tuned to enhance its efficacy in our designated tasks. This approach is in line with cutting-edge methodologies in natural language processing. For evaluation, we select up to 1,000 instances from three distinct random splits, ensuring a diverse and representative test dataset. This strategy aims to emulate real-world scenarios and offers a comprehensive evaluation of the model’s capabilities.
4.2 Compared models
We evaluate the performance of several prominent Large language models. ChatGPT (gpt-3.5-turbo-0301) is renowned for its conversational abilities and versatility in various language tasks, initially based on the GPT-2 model from Radford et al. (2019). Vicuna, credited to Chiang et al. (2023), stands out for its efficiency, being fine-tuned on a LLaMA base model using conversational data. Similarly, Alpaca, developed by Taori et al. (2023), uses the LLaMA architecture but is specifically tuned for domain-specific dialogues. Our focus, Llama 2, is fine-tuned following tou (2023) guidelines, showcasing its capabilities in complex language processing scenarios and offering insights into the evolving landscape of vaious NLP tasks.
4.3 LoRA approach for fine-tuning
Low-Rank Adapters (LoRA) represent a pivotal advancement in the field of natural language processing, particularly in the fine-tuning of large language models. First introduced by Hu et al. (2021), LoRA allows for the efficient and effective modification of pre-trained models, significantly reducing the computational overhead typically associated with training large-scale neural networks. In our implementation, we integrate LoRA adapters directly into the attention mechanisms and train both the language model (LM) heads and the adapters concurrently, diverging from traditional methods that prioritize the initial training of embeddings alone, as seen in models like the Chinese LLaMA Cui et al. (2023). Our infrastructure utilizes the robust capabilities of six NVIDIA A100 GPUs, each with 46GB of memory, enabling us to train our model, DevAssistLlama, over three epochs spanning a total of 63 hours. We showcased important hyperparameters in Table 6.
5 Results
@@ 5.1 Evaluation metrics
Learning Rate: 2e-4| Max sequence length: 4096|
LoRA Alpha: 16| Training Precision: FP16|
Epochs: 3| Optim: paged_adamw_32bit|
fp16=True| bf16=True| Warmup ratio=0.03|
LR scheduler=Cosine| Trainer= SFTT|
@@
Table 6: Summary of important hyperparameters
6 Limitation and future work